# Boosting Mixed-Initiative Co-Creativity in Game Design: A Tutorial

### v1

@@ Abstract.
copyright: noneccs: Applied computing Computer gamesccs: Human-centered computing Collaborative interactionccs: General and reference Surveys and overviews
@@ 1. Introduction
Mixed-initiative and co-creativity are two increasingly prominent subjects within the research areas of human-computer interaction and computational creativity. As AI systems are becoming increasingly powerful and sophisticated, we are shifting from the view of AI as a tool towards AI as an active participant. Thus, there is a growing concern among researchers in exploring how we can benefit from human-AI partnerships. Particularly, there has been extensive research into the interaction between humans and computational systems to foster a collaborative environment where both agents exhibit initiative and proactivity, hopefully enhancing the creative process (Yannakakis et al., 2014). As stated by Davis(Davis, 2013), “human-computer co-creativity describes a situation in which the human and the computer improvise in real-time to generate a creative product”. Mixed-initiative systems were defined in 1970 by Carbonell (Carbornell, 1970) as those “in which both humans and machines can make contributions to a problem solution, often without being asked explicitly”. More recently, Yannakakis et al. (Yannakakis et al., 2014) introduced the term mixed-initiative co-creation, describing it as “the task of creating artifacts via the interaction of a human initiative and a computational initiative”, where they emphasize the proactivity of both parties. Although the use of different terms to describe mixed-initiative may cause some ambiguity, the addition of co-creativity, or an alternative term such as human-computer co-creativity, can serve as a highlight of the close collaboration among the agents and the creative nature of the process and its resulting output. Throughout this paper, we refer to mixed-initiative co-creativity (MI-CCy) as we are focusing on a strong creative partnership between humans and computers.
@@
Video games are a domain of particular interest for the application of mixed-initiative co-creative approaches. As the market grows and the players’ expectations increase, so does the time, money, and effort needed to create games with large amounts of unique content that meet the expected quality (Shaker et al., 2016; Baldwin et al., 2017a; Walton et al., 2022). For this reason, several support tools (e.g. level editors or game engines) have been developed to streamline game development processes (Yannakakis et al., 2014). Along with these, there has been major research and development of tools to automate content generation, known as procedural content generation (PCG) (Shaker et al., 2016; Baldwin et al., 2017a). While very beneficial, there is no human-computer co-creation in the use of these tools. When using support frameworks, the creative process relies merely on human initiative. As for PCG, the human creative input is very limited, mainly consisting of the choice of parameters before and/or after the generative process (Yannakakis et al., 2014). The designer’s lack of control often results in a lower quality of user experience, with generated content coming off as uninteresting and predictable (Karavolos et al., 2015; Shaker et al., 2016). Mixed-initiative approaches applied to game design have the potential to address these issues, as humans and machines work together, both proactively contributing to the creative process.
@@ 2. Classes of Game Content
The game space represents the environment in which the game unfolds. It can range from a preliminary outline of the spatial structure, the components it may contain, and their respective placements, to a high-level representation, complete with the final look and all the associated game bits in place. When talking about game space, we mainly refer to maps, which can be indoor or outdoor and are composed of navigable space, obstacles (e.g., walls, rocks, and rivers), and interactive elements, such as collectibles. Maps may also have special features such as teleportation areas. Some common map types include dungeons, consisting of a group of rooms interconnected by corridors containing enemies and treasures (Dahlskog et al., 2015), and platform games, which usually feature suspended platforms with varying shapes and sizes placed at different heights.
@@
In this tutorial, we define the game behaviors class as the predefined way game elements such as objects and characters act and interact with their surroundings. This includes their relationship with the physics of the game (e.g., whether they float or react to gravity) and their behavior in different kinds of interactions such as moving, expanding, exploding, and disintegrating. It also includes the various mechanics of each game element and their specific characteristics (e.g., jumping, running, throwing, and sliding). The characters’ own way of speaking and acting and their personality traits are also associated with game behaviors.
@@
Game systems combine sets of rules and algorithms to provide and manage complex (and often sustainable) interactions between game elements at a large scale. Hendrikx et al. identify four main types of game systems: ecosystems, which dictate the evolution and interplay of plant and animal life; road networks, which affect traffic, resource exchanges, and access to services in outdoor locations; urban environments, concerning the efficient organization of clusters of buildings for the resident population; and entity behavior, such as complex interactions of non-playable characters with the player and as a group (Hendrikx et al., 2013).
@@
Game scenarios outline the sequence of game events and how they unfold (Hendrikx et al., 2013). In earlier stages of game design, game scenarios can be portrayed through media like storyboards or diagrams. The story of a game is a prime example of a game scenario, primarily prominent in narrative-driven games. Puzzles provide a space of possible solutions that may require the occurrence of very specific orders of events. Moreover, this class encompasses the strategic organization of game levels and the careful planning of level progression in accordance with an underlying mission or set of objectives.
@@
Finally, we must also refer to the most general and conceptual aspects of a game such as rules, goals, genre, aesthetic theme, historical context, depicted reality/universe, and narrative tone. Converting these concepts and ideas into a game involves exploring the various classes of content mentioned above. In fact, a game can simply be designed in our head (Schell, 2008), but it is by going through the myriad steps of game content creation that we can turn it into a physically playable artifact.
@@ 3. Mixed-Initiative in Game Design: Literature Review
This section gathers a selection of current works involving the development and exploration of mixed-initiative approaches for game design processes. The selection made was based on the search for the keywords “mixed-initiative” along with “game design”, “game” or “games” in Google Scholar, focusing on relevance and topicality. The same search combinations were made by replacing “mixed-initiative” with “human-computer co-creativity”, “human-computer co-creation”, “human-AI co-creativity”, “human-AI co-creation” and “co-creative AI” to ensure that no work would be missed due to the use of different terms to express the same idea. To keep a focus on video games, the selection of works we conducted prioritized mixed-initiative approaches exclusively applied to game design contexts. While we did come across some studies that could potentially have application in game design but did not primarily target it, such as mixed-initiative approaches for writing poetry (Oliveira et al., 2017; Chakrabarty et al., 2022) or creating narratives (Samuel et al., 2016; Stefnisson and Thue, 2018; Kreminski et al., 2022; Yuan et al., 2022; Lee et al., 2022), we chose not to include them as our case studies.
@@
We decided to narrow down our selection of works to ten candidates in order to have a sample size that is sufficiently small to allow for a detailed analysis yet large enough to be representative of the state of the art. We aimed to include at least one example for each class of game content. For those classes that are more prevalent in MI-CCy systems, we included more than one exemplary work, reflecting the prevailing trends in the field. These classes are game space, with a particular focus on map design, followed by game scenarios, which places greater emphasis on narrative design, and then game bits. The game systems class does not have any work selected to be analyzed with our framework due to the lack of literature that sufficiently meets our MI-CCy criteria. In determining which works to include for each game content class, our primary criterion was to feature recent works, as well as works that held significant relevance in this research area. We sought to ensure there were substantial differences among the various candidates to guarantee diversity in our study. Another pivotal criterion in the selection process was the availability of the system online, either through a web application or repository, for testing purposes. The direct interaction with the systems allowed us to conduct a more in-depth analysis. The works are presented below, grouped according to the class of game content most prevalently addressed by each approach.
@@ 3.1. Game Bits
Tools specifically aimed at creating game bits based on a mixed-initiative co-creative approach are scarce in the literature. There are several co-creative systems for art creation; however, that can easily be used to produce game content. This is the case of systems for co-creative sketching, such as Creative Sketching Partner (Karimi et al., 2020), where a user starts drawing and can ask the artificial agent for inspiration at any time. The system responds with a sketch with a desired amount of visual and conceptual similarity to the user’s sketch. Then, the user may be inspired to make changes to their sketch and repeat the process if they wish to. Despite the diversity of mixed-initiative tools potentially applicable to game design, to stay on subject, our literature selection only admits works that specifically target game content creation. An initial endeavor in mixed-initiative targeted at generating game bits is the work of Liapis et al. (Liapis et al., 2012a, b) They introduced an approach for generating personalized content based on user preference regarding visual aesthetics. The framework presents eight spaceships that were optimized offline, from which the user selects their preferred one. The system then generates a new set of spaceships, enhancing the visual features behind the user’s choice via an adaptive aesthetic model. After a few iterations, the newly evolved content is more likely to be appealing to the user.
@@ 3.1.1. AI Spaceship Generator
Gallotta et al. (Gallotta et al., 2023) presented an approach for content generation aligned with the user’s preferences by incorporating preference learning into quality-diversity algorithms. To showcase this approach, they implemented an application to generate 3D spaceships for Space Engineers, a game that allows players to construct block-based structures. The interface of the AI Spaceship Generator (Fig. 1) presents a grid with an initial population of spaceship suggestions. The user can select any of these suggestions and see its 3D preview along with a list of its in-game properties. If the user appreciates the selected spaceship, they can press the “Evolve from Random Spaceship” button and the system will generate more suggestions via a predictive model of the preferences of the human designer. Otherwise, the user can press the “Evolve from Random Spaceship” button to be presented with more random suggestions. As the spaceship population is updated, the user can select a new spaceship and repeat the process. With each generation based on user selection, the system continues to learn preferences for the newly evolved content to stay aligned with the user’s intent. The user can also change the color of the spaceship and a few generation settings. The spaceship resulting from the iterative process can be downloaded and imported into Space Engineers.
Extending procedural content generation systems to mixed-initiative settings by solely accounting for user selection is a common but limited approach if the aim is to afford human-computer co-creativity. We discuss this further in Subsection 4.2.
Figure 1. Interface of the AI Spaceship Generator (Gallotta et al., 2023). Top left - the spaceship population; top middle - 3D preview of the selected spaceship; top right - a list of the spaceship properties; bottom left - controls to generate more spaceships or reset the population; bottom middle - option to change the spaceship color and option to download its blueprint; bottom right - application logs. (Available at https://github.com/arayabrain/space-engineers-ai-spaceship-generator)
@@ 3.1.2. Sprite Editor for Pixel Art Characters
Coutinho and Chaimowicz (dos Santos Coutinho and Chaimowicz, 2023) introduced an approach for creating sprites for pixel art characters facing four directions (front, back, left, and right). To this end, they developed a sprite editor (Fig. 2) in which the human agent can interact with the computational agent to generate the various poses. The system contains a main canvas for the user to view and edit their selected pose. The human agent initiates the collaborative process by starting to draw the character on the main canvas. At any time, the designer can drag the pose they drew into a suggestion slot for a different pose (e.g., dragging the front view into the right view suggestion slot). A trained generative model takes this image as input and generates an image for the target pose, displaying it in the suggestion slot. The human agent can engage in multiple iterative steps, either directly editing any pose view or asking the computational agent to generate a new suggestion from any view that is not the same direction as the target.
Figure 2. Interface of the sprite editor (dos Santos Coutinho and Chaimowicz, 2023). Left - the toolbar; middle - the main canvas displaying the selected pose view; right - a panel with the four pose views and a panel with the suggestion slots for each pose view on its left side. (Available at https://fegemo.github.io/sprite-editor-gm/)
@@ 3.2.1. Sentient Sketchbook
Liapis et al. (Liapis et al., 2013; Yannakakis et al., 2014) created a tool for the design of game levels. The Sentient Sketchbook (Fig. 3) enables users to create and edit levels using map sketches as simplified representations of levels. As the human designer engages with the tool through sketching, the system dynamically provides suggestions that are evolved variations of their current sketch. These suggested alternative map designs are generated using a genetic search algorithm that aims to optimize the levels’ fitness scores or produce visually diverse maps. The tools’ interface displays the fitness scores and other metrics of the current sketch for the user to compare it to the scores and metrics of each suggestion. The human agent can choose to replace their current sketch with a suggested alternative at any time. Once satisfied with the outcome, the designer can request the system to convert the map sketch into a playable level with enhanced detail.
Figure 3. The Sentient Sketchbook tool (Yannakakis et al., 2014). Left - the sketch editor; middle - the tile palette, the map display menu, and an overview of the map’s fitness dimensions and metrics; right - the automatically generated map suggestions. (Available at https://www.sentientsketchbook.com/download.php)

### v2

@@ Abstract.
copyright: ccjournal: CSURjournalyear: 2025journalvolume: 1article: 1publicationmonth: 8doi: 10.1145/3759914ccs: Applied computing Computer gamesccs: Human-centered computing Collaborative interactionccs: General and reference Surveys and overviews
@@ 1. Introduction
Mixed-initiative and co-creativity are two increasingly prominent subjects within the research areas of human-computer interaction and computational creativity. As AI systems are becoming increasingly powerful and sophisticated, we are shifting from the view of AI as a tool towards AI as an active participant. Thus, there is a growing concern among researchers in exploring how we can benefit from human-AI partnerships. Particularly, there has been extensive research into the interaction between humans and computational systems to foster a collaborative environment where both agents exhibit initiative and proactivity, hopefully enhancing the creative process (Yannakakis et al., 2014). As stated by Davis(Davis, 2013), “human-computer co-creativity describes a situation in which the human and the computer improvise in real-time to generate a creative product”. Mixed-initiative systems were defined by Carbonell (Carbornell, 1970) as those “in which both humans and machines can make contributions to a problem solution, often without being asked explicitly”. More recently, Yannakakis et al. (Yannakakis et al., 2014) introduced the term mixed-initiative co-creation, describing it as “the task of creating artifacts via the interaction of a human initiative and a computational initiative”, emphasizing the proactivity of both parties. Although the use of different terms to describe mixed-initiative may cause some ambiguity, the addition of co-creativity, or an alternative term such as human-computer co-creativity, can serve as a highlight of the close collaboration among the agents and the creative nature of the process and its resulting output. Throughout this paper, we refer to mixed-initiative co-creativity (MI-CCy) as we are focusing on a strong creative partnership between humans and computers.
@@
Video games are a domain of particular interest for the application of mixed-initiative co-creative approaches. As the market grows and the players’ expectations increase, so does the time, money, and effort needed to create games with large amounts of unique content that meet the expected quality (Shaker et al., 2016; Baldwin et al., 2017a; Walton et al., 2022). For this reason, several support tools (e.g., level editors or game engines) have been developed to streamline game development processes (Yannakakis et al., 2014). Along with these, there has been major research and development of tools to automate content generation, known as procedural content generation (PCG) (Shaker et al., 2016; Baldwin et al., 2017a). While very beneficial, there is no human-computer co-creation in the use of these tools. When using support frameworks, the creative process relies merely on human initiative. As for PCG, the human creative input is very limited, mainly consisting of the choice of parameters before and/or after the generative process (Yannakakis et al., 2014). The designer’s lack of control often results in a lower quality of user experience, with generated content coming off as uninteresting and predictable (Karavolos et al., 2015; Shaker et al., 2016). Mixed-initiative approaches applied to game design have the potential to address these issues, as humans and machines work together, both proactively contributing to the creative process.
@@ 2. Classes of Game Content
The game space represents the environment in which the game unfolds. It can range from a preliminary outline of the spatial structure, the components it may contain, and their respective placements, to a high-level representation, complete with the final look and all the associated game bits in place. When talking about game space, we mainly refer to maps, which can be indoor or outdoor and are composed of navigable space, obstacles (e.g., walls, rocks, and rivers), and interactive elements, such as collectibles. Some common map types include dungeons, consisting of a group of rooms interconnected by corridors containing enemies and treasures (Dahlskog et al., 2015), and platform games, which usually feature suspended platforms with varying shapes and sizes placed at different heights.
@@
In this tutorial, we define the game behaviors class as the predefined way game elements, such as objects and characters, act and interact with their surroundings. This includes their relationship with the physics of the game (e.g., whether they float or react to gravity) and their behavior in different kinds of interactions such as moving, expanding, exploding, and disintegrating. It also includes the various mechanics of each game element and their specific characteristics (e.g., jumping, running, throwing, and sliding). The characters’ own way of speaking and acting and their personality traits are also associated with game behaviors.
@@
Game systems combine sets of rules and algorithms to provide and manage complex (and often sustainable) interactions between game elements at a large scale. Hendrikx et al. (Hendrikx et al., 2013) identify four main types of game systems: ecosystems, which dictate the evolution and interplay of plant and animal life; road networks, which affect traffic, resource exchanges, and access to services in outdoor locations; urban environments, concerning the efficient organization of clusters of buildings for the resident population; and entity behavior, such as complex interactions of non-playable characters with the player and as a group.
@@
Game scenarios outline the sequence of game events and how they unfold (Hendrikx et al., 2013). In earlier stages of game design, game scenarios can be portrayed through media such as storyboards or diagrams. The story of a game is a prime example of a game scenario, primarily prominent in narrative-driven games. Puzzles provide a space of possible solutions that may require the occurrence of very specific orders of events. Moreover, this class encompasses the strategic organization of game levels and the careful planning of level progression in accordance with an underlying mission or set of objectives.
@@
Finally, we must refer to the most general and conceptual aspects of a game, such as rules, goals, genre, aesthetic theme, historical context, depicted universe, and narrative tone. Converting these concepts and ideas into a game involves exploring the various classes of content mentioned above. In fact, a game can be designed in our head (Schell, 2008), but by going through the myriad steps of creating game content, we can turn it into a physically playable artifact.
@@ 3. Mixed-Initiative in Game Design: Literature Review
This section gathers a selection of current works involving the development and exploration of mixed-initiative co-creative approaches for game design processes. The selection made was based on the search for the keywords “mixed-initiative” along with “game design”, “game” or “games” in Google Scholar, focusing on relevance and topicality. The same search combinations were made by replacing “mixed-initiative” with “human-computer co-creativity”, “human-computer co-creation”, “human-AI co-creativity”, “human-AI co-creation”, and “co-creative AI” so that no work would be missed due to the use of different terms to express the same idea. The focus on these terminologies ensures that the selected tools not only involve the use of AI for creative tasks in game design but also provide a clear description of how human and computational agents interact and collaborate creatively.
To keep the focus on games, the selection of works we conducted prioritized mixed-initiative approaches exclusively applied to game design contexts. Although we did find some studies that could potentially have application in game design but did not primarily target it, such as mixed-initiative approaches to writing poetry (Gonçalo Oliveira et al., 2017; Chakrabarty et al., 2022) or creating narratives (Samuel et al., 2016; Stefnisson and Thue, 2018; Kreminski et al., 2022; Yuan et al., 2022; Lee et al., 2022), we chose not to include them in our case studies, as they are not tailored to the specific needs and structures of game development. Additionally, our selection focuses exclusively on video games and does not include related media such as board games or tabletop role-playing games. While AI applications in these domains are emerging (Triyason, 2023), especially through general-purpose technologies like ChatGPT, systems specifically designed to support mixed-initiative co-creativity in these contexts are still scarce. Our focus on video games reflects both the current state of the art and our interest in purpose-built game design tools.
@@
We decided to narrow down our selection of works to ten candidates in order to have a sample size that is sufficiently small to allow for a detailed analysis yet large enough to be representative of the state of the art. We aim to include at least one example for each class of game content. For those classes that are more prevalent in MI-CCy systems, we included more than one exemplary work, reflecting the prevailing trends in the field. These classes are game space, with a particular focus on map design, followed by game scenarios, which places greater emphasis on narrative design, and then game bits. The game systems class does not have any work selected to be analyzed with our framework due to the lack of literature that sufficiently meets our MI-CCy criteria. In determining which works to include for each game content class, our primary criterion was to feature recent works, as well as works that held significant relevance in this research area. We sought to ensure that there were substantial differences among the various candidates to guarantee diversity in our study. Another crucial criterion in the selection process was the availability of the system online, through a web application or repository, for testing purposes. The direct interaction with the systems allowed us to conduct a more in-depth analysis.
It is worth noting that our selection does not include systems from the commercial sector. This gap is due to two key factors: first, the lack of public and free access to most commercial tools, which prevents direct evaluation; and second, our aim to maintain a focused scope on tools that explicitly support mixed-initiative co-creativity in the context of game design. While there are notable commercial tools that apply AI to creative processes, to the best of our knowledge, most are either not specifically tailored for game design (e.g., KREA (Rodriguez Prado et al., 2024)) or do not meet the criteria for MI-CCy or provide sufficient detail about the human–AI collaboration to be meaningfully analyzed within our framework (e.g., Popul8 (Dias et al., 2024)). That said, we recognize that commercial developments in this area are highly relevant and contribute significantly to advancing AI-supported creativity in practice. The final candidates are presented below, grouped according to the class of game content most prevalently addressed by each approach.
@@ 3.1. Game Bits
Mixed-initiative tools specifically aimed at creating game bits are scarce in the literature. There are several co-creative systems for art creation; however, they can easily be used to produce game content. This is the case of systems for co-creative sketching, such as Creative Sketching Partner (Karimi et al., 2020), where a user starts drawing and can ask the artificial agent for inspiration at any time. The system responds with a sketch with a desired amount of visual and conceptual similarity to the user’s sketch. Then, the user may be inspired to make changes to their sketch and repeat the process if they wish to. Despite the diversity of mixed-initiative tools potentially applicable to game design, to stay on subject, our literature selection only admits works that specifically target game content creation. An initial endeavor in mixed-initiative targeted at generating game bits is the work of Liapis et al. (Liapis et al., 2012a, b). They introduced an approach for generating personalized content based on user preference regarding visual aesthetics. The framework presents eight spaceships that were optimized offline, from which the user selects their preferred one. The system then generates a new set of spaceships, enhancing the visual features behind the user’s choice via an adaptive aesthetic model. After a few iterations, the newly evolved content is more likely to be appealing to the user.
@@ 3.1.1. AI Spaceship Generator
Gallotta et al. (Gallotta et al., 2023) presented an approach for content generation aligned with the user’s preferences by incorporating preference learning into quality-diversity algorithms. To showcase this approach, they implemented an application to generate 3D spaceships for Space Engineers, a game that allows players to construct block-based structures. The interface of the AI Spaceship Generator (Fig. 1) presents a grid with an initial population of spaceship suggestions. The user can select any of these suggestions and see its 3D preview along with a list of its in-game properties. If the user appreciates the selected spaceship, they can press the “Evolve from Selected Spaceship” button and the system will generate more suggestions via a predictive model of the preferences of the human designer. Otherwise, the user can press the “Evolve from Random Spaceship” button to be presented with more random suggestions. As the spaceship population is updated, the user can select a new spaceship and repeat the process. With each generation based on user selection, the system continues to learn preferences for the newly evolved content to stay aligned with the user’s intent. The user can also change the color of the spaceship and a few generation settings. The spaceship resulting from the iterative process can be downloaded and imported into Space Engineers. Extending procedural content generation systems to mixed-initiative settings by solely accounting for user selection is a common but limited approach if the aim is to afford human-computer co-creativity. We discuss this further in Subsection 4.2.
Figure 1. Interface of the AI Spaceship Generator (Gallotta et al., 2023). Top left — the spaceship population; top middle — 3D preview of the selected spaceship; top right — a list of the spaceship properties; bottom left — controls to generate more spaceships or reset the population; bottom middle - option to change the spaceship color and option to download its blueprint; bottom right - application logs. (Available at https://github.com/arayabrain/space-engineers-ai-spaceship-generator)
@@ 3.1.2. Sprite Editor for Pixel Art Characters
Coutinho and Chaimowicz (Coutinho and Chaimowicz, 2024) introduced an approach for creating sprites for pixel art characters facing four directions (front, back, left, and right). To this end, they developed a sprite editor (Fig. 2) in which the human agent can interact with the computational agent to generate the various poses. The system contains a main canvas for the user to view and edit their selected pose. The human agent initiates the collaborative process by starting to draw the character on the main canvas. At any time, the designer can drag the pose they drew into a suggestion slot for a different pose (e.g., dragging the front view into the right view suggestion slot). A trained generative model takes this image as input and generates an image for the target pose, displaying it in the suggestion slot. The human agent can engage in multiple iterative steps, either directly editing any pose view or asking the computational agent to generate a new suggestion from any view that is not the same direction as the target.
Figure 2. Interface of the sprite editor (Coutinho and Chaimowicz, 2024). Left — the toolbar; middle — the main canvas displaying the selected pose view; right — a panel with the four pose views and a panel with the suggestion slots for each pose view on its left side. (Available at https://fegemo.github.io/sprite-editor-gm/)
@@ 3.2.1. Sentient Sketchbook
Liapis et al. (Liapis et al., 2013; Yannakakis et al., 2014) created a tool for the design of game levels. The Sentient Sketchbook (Fig. 3) enables users to create and edit levels using map sketches as simplified representations of levels. As the human designer engages with the tool through sketching, the system dynamically provides suggestions that are evolved variations of their current sketch. These suggested alternative map designs are generated using a genetic search algorithm that aims to optimize the levels’ fitness scores or produce visually diverse maps. The tools’ interface displays the fitness scores and other metrics of the current sketch for the user to compare them to the scores and metrics of each suggestion. The human agent can choose to replace their current sketch with a suggested alternative at any time. Once satisfied with the outcome, the designer can request the system to convert the map sketch into a playable 2D or 3D level with enhanced detail.
Figure 3. The Sentient Sketchbook tool (Yannakakis et al., 2014). Left — the sketch editor; middle — the tile palette, the map display menu, and an overview of the map’s fitness dimensions and metrics; right — the automatically generated map suggestions. (Available at https://www.sentientsketchbook.com/download.php)